OpenStudy (anonymous):

1 7 11 -- + -- = -- 9a 3a 18

OpenStudy (kc_kennylau):

Multiply the second fraction's numerator and denominator by 3.

OpenStudy (anonymous):

1 21 11 -- + --- = -- 9a 9a 18

OpenStudy (kc_kennylau):

Then add the first fraction and the second fraction together.

OpenStudy (anonymous):

22 11 -- = -- 9a 18

OpenStudy (kc_kennylau):

Cross multiplication \(\left(\dfrac ab=\dfrac cd\Rightarrow ad=bc\right)\)

OpenStudy (anonymous):

i got 4 as my final answer

OpenStudy (kc_kennylau):

yes, that is correct.

OpenStudy (anonymous):

thanks so much!!!

OpenStudy (anonymous):

way to go

OpenStudy (kc_kennylau):

no problem :)